find the second derivative of x^1/2
1
Expert's answer
2012-10-31T11:40:42-0400
(x^1/2)'=1/2*(x^(-1/2))
(x^1/2)''=(1/2*(x^(-1/2)))'=1/2*(-1/2)*(x^(-3/2))=-1/4*x^(-3/2)